OpenStudy (anonymous):

If a wave has amplitude of 2 meters, a wavelength of 2 meters, and a frequency of 10 Hz, and a period of 1 second, then at what speed is the wave moving? A. 40 m/s B. 20 m/s C. 10 m/s D. 5 m/s

OpenStudy (anonymous):

frequency=1/T it frequency =10 then time =1/10 if one cycle takes T=1/10 distance traveled=2m then for time 1 sec distance traveled = 2*10m=20m speed 20m/sec

OpenStudy (anonymous):

@KendallLittell

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

im doing this summer homework on line and i have like two more units to go you guys are really helping me !

OpenStudy (anonymous):

well try to understand how to do rather than just finishing homework wil help you alot in future take care